### v1

# Bridging the Gap: OPTICAM Reveals the Hidden Spin of the WZ Sge Star GOTO 065054.49+593624.51

Abstract
WZ Sge stars are highly evolved accreting white dwarf systems (AWDs) exhibiting remarkably large amplitude outbursts (a.k.a. super-outbursts), typically followed by short rebrightenings/echo outbursts. These systems have some of the lowest mass transfer rates among AWDs, making even low magnetic fields dynamically important. Such magnetic fields are often invoked to explain the phenomenology observed in these systems, such as their X-ray luminosity and long periods of quiescence (30+ years). However, the detection of these is very elusive given the quenching of the accretion columns during outburst and the low luminosity of these systems during quiescence. Here we present high-cadence multi-band observations with OPTICAM of the recent outburst of the recently discovered WZ Sge star GOTO065054.49+593624.51, during the end of the main outburst and the dip in-between rebrightenings, covering 2 orders of magnitude in brightness. Our observations reveal the presence of a statistically significant signal with $P_{\omega}\simeq 148$ seconds in the bluer ($g$) band which is detected only during the dip between the main outburst and the rebrigthenings. We interpret this signal as the spin period of the AWD. If confirmed, GOTO 0650 would bridge the gap between intermediate- and fast-rotating intermediate polars (IPs) below the period gap.
keywords:
accretion, accretion discs – stars: dwarf novae – stars: individual: GOTO065054.49+593624.51
pubyear: 2025pagerange: Bridging the Gap: OPTICAM Reveals the Hidden Spin of the WZ Sge Star GOTO 065054.49+593624.51Bridging the Gap: OPTICAM Reveals the Hidden Spin of the WZ Sge Star GOTO 065054.49+593624.51
1 Introduction
Cataclysmic variable stars (CVs) are binary systems in which a low-mass secondary star, filling its Roche lobe, transfers mass to a white dwarf (WD) primary. In systems where the magnetic field is dynamically unimportant, the accretion process is mediated by an accretion disk. If the magnetic field of the WD is sufficient to prevent the formation of the disk, this gives rise to the so-called polars. If, on the other hand, a disc can form but is truncated by the magnetosphere of the central WD, the systems are classified as intermediate-polars (IPs; see e.g. Warner, 1995, for a review). Disk-mediated systems with mass transfer rates sufficiently low to prevent the disk from remaining persistently in a fully ionized state are classified as Dwarf novae (DNe). DNe are characterized by exhibiting eruptions where the system brightens typically 2-5 magnitudes for a relatively short period of time. Such eruptions are thought to be the consequence of a sudden increase in the mass accretion rate onto the compact object due to the accretion disk transitioning from a neutral to a fully ionized state (e.g., Lasota, 2016).
WZ Sge-systems are highly-evolved short orbital period DNe which exhibit even larger amplitude outbursts (up to $\sim 8-9$ mag). These so-called super-outbursts are characterized by prolonged outburst durations (approximately 30 days), and the appearance of superhumps — low-amplitude variations near the system’s orbital period — shortly after reaching peak brightness (Smak, 1993, 2009; Patterson et al., 2005; Kato, 2015; Hameury, 2020). These super-outbursts are typically followed by a series of short-term rebrightenings or echo outbursts which seem to be characteristic of accreting binaries with extreme mass-ratios. These rebrightenings cannot be explained by the standard disk instability model (DIM; e.g. Hameury, 2020). To reconcile observations with theoretical predictions, Campana et al. (2018) recently proposed an alternative explanation where the luminosity drops during outburst decay signal a temporary transition to a magnetic propeller state. This is an appealing framework as it could help to explain the behaviour of accreting systems at different scales, i.e. X-ray binaries, young stellar objects and CVs. This framework also aligns well with the proposed spin period of 27.87 seconds of WZ Sge (Patterson, 1980; Patterson et al., 1998). However, using HST time-resolved spectroscopy, Georganti et al. (2022) showed that WZ Sge does not exhibit any propeller footprint during the proposed propeller state (c.f. Eracleous & Horne, 1996). Nonetheless, truncation of the inner disk is required to explain the relatively high X-ray luminosity. Against this background, Hameury (2020) proposed that the X-ray excess observed in WZ Sge could (also) be a consequence of evaporation of the inner disk, while the quiescent phases between outbursts are produced by overflows of the mass-transfer stream impacting the inner regions of the disc, similar to what is proposed to explain IW And-systems and some nova-like CVs (Kimura et al., 2020; Castro Segura et al., 2021). Regardless of whether magnetic fields are responsible for the X-ray luminosities in WZ Sge stars, they are often invoked to explain the very long recurrence timescales in these systems (Kato, 2015).
In this Letter we report high cadence multi-band photometry of the newly discovered transient GOTO065054.49+593624.51 (hereafter GOTO 0650) obtained with OPTICAM (Castro et al., 2019, 2024). GOTO 0650 was discovered by the Gravitational-wave Optical Transient Observer (GOTO; Steeghs et al., 2022; Dyer et al., 2024) on Oct. 4 2024 03:36:36 UT (Killestein et al., 2024a). The transient reached an L band ($\sim g+r$) magnitude of 13.7 in the discovery images and was associated with a $\sim 22^{nd}$ mag quiescent counterpart implying an outburst amplitude of $\sim 8.5$ mag, typical of WZ Sge stars. The nature of the transient was confirmed by spectroscopic observations and ultraviolet photometry (e.g. Killestein et al., 2024b; Bhattacharya & Bhattacharyya, 2024). Using data from the American Association of Variable Star Observers (AAVSO) and the Zwicky Transient Facility (ZTF; Bellm et al., 2019a, b), we show in Figure 1 how the outburst evolution and rebrigthenings are consistent with the classification above. Tampo (2024) reported the onset of regular super-humps after a type-E recovery of a dip $\sim 15\,{\rm d}$ from the beginning of the outburst (e.g. Kimura et al., 2018), and suggested GOTO 0650 is a period bouncer with an orbital period close to the superhump period (${\rm P_{sh}}=91.05\pm 3$ min), again, in line with GOTO 0650 being a member of the WZ Sge-type DNe.
Figure 1: Outburst evolution of GOTO 0650 as reported by the AAVSO (circles) and ZTF (diamonds for psf- and stars for forced-photometry). The different colours indicate the magnitude in different bands. The coloured horizontal bands encompass the ±$\pm 1\sigma$ error around the weighted mean of the quiescent photometry obtained with forced photometry on ZTF during 9-25 days before the onset of the superoutburst, the precise values are $g=22.25\pm 0.03$ mag and $r=22.351\pm 0.005$ mag. The arrows highlight the maxima of the post-outburst re-brightening/-flares. The time of the high-cadence observations with OPTICAM are indicated with the grey shaded regions, the epoch ID for each observation reported in this paper is indicated on the top axis. (*) denotes the epoch with a statistically significant detection of the $148$s signal.
2 Observations and data reduction
The OPtical TIming CAMera (OPTICAM; Castro et al., 2019, 2024) is a new high-cadence, multi-band camera mounted on the 2.1-meter telescope at the San Pedro Mártir Observatory (OAN-SPM), in México. OPTICAM is equipped with three Andor Zyla 4.2-Plus sCMOS cameras with three $2048\times 2048$ pixels and a set of SDSS filters ($u^{\prime}g^{\prime}r^{\prime}i^{\prime}z^{\prime}$), allowing coverage in the $3200{\rm AA}-1.1{\rm\mu m}$ range. The field of view (FoV), is $\approx 5\times 5~{}{\rm arcmin^{2}}$ with a pixel-scale of $\simeq 0.15^{\prime\prime}/{\rm pix}$.
GOTO 0650 was observed in the $g^{\prime},r^{\prime},i^{\prime}$ bands during the nights of October 26, 28, 30 and November 2nd and 4th of 2024, with a cadence of $3$-s, $3$-s, $5$-s, $15$-s, $15$-s respectively. The typical length of the observations was $\simeq 4.2$ hrs; the details of the observations can be found in Table 1. The data were primarily reduced using version 1.13.0 of the photutils Python package (Bradley et al., 2024). Cosmic rays were clipped from all images using the L.A.Cosmic algorithm (van Dokkum, 2001) as implemented in Astro-SCRAPPY version 1.2.0 (McCully et al., 2018). For each image, two dimensional background images were calculated using the SExtractorBackground estimator from photutils with a box_size of 32 pixels. Image segmentation was used to identify sources in the background and subtracted images using the SourceFinder routine from photutils, which combines source detection and deblending. To identify sources, we set the npixels parameter to 8 pixels, and the threshold parameter to $5\times{\rm background~{}RMS}$. The background RMS was estimated using StdBackgroundRMS. All other parameters were left to their default values.
We used the optimal photometry algorithm described in Naylor (1998) to perform photometry on the background-subtracted images. We model the PSF of OPTICAM’s three cameras using two-dimensional Gaussians for each camera. We stacked the images from each epoch for each camera to get the PSF parameters. We then identified sources in the stacked images using SourceFinder, which assigns a semi-major and semi-minor standard deviation to each source. The PSF for each camera in each epoch was then modelled using the median semi-major and semi-minor standard deviations from the corresponding stacked image. We then used differential photometry to correct for atmospheric variability; relative light curves were computed using the same (non-variable) reference stars in each night.
Table 1: Observing log. The star in the epoch number denotes statistically significant detection of short-term periodicity (see Sec. 3 for details).
Epoch MJDstart filters exposure time obs. length binning
day seconds hours
$1$ $60609.31$ $gri$ $3$ $4.65$ $2\times 2$
$2$ $60611.33$ $gri$ $3$ $4.54$ $2\times 2$
$3$ $60613.38$ $gri$ $5$ $3.62$ $2\times 2$
$4^{*}$ $60616.39$ $gri$ $15$ $3.34$ $2\times 2$
$5$ $60618.29$ $gri$ $15$ $5.89$ $2\times 2$
3 Analysis and results
The high cadence photometric light curves are presented in the left panels of Figures 2 and 3, the comparison stars in these figures are always the same, so the change in relative flux between epochs is meaningful. As can be seen in Figure 1, epochs 1 and 2 were taken towards the end of the superoutburst, epoch 3 was taken during the decline, epochs 4 and 5 occurred during the onset of the dip between the main outburst structure and the rebrightenings, when the source was $\sim 4-5$ mags fainter than the first observation but $\sim 3$ mags brighter than the quiescent level. During epochs 1–3, the source was dominated by the regular superhumps reported by Tampo (2024), with the dispersion in the light curves reducing as the accretion disk got fainter. In epochs 4 and 5 there is still a slow modulation present with similar frequency as before, as can be seen in Figure 3, the general trend is similar but the $g$-band exhibits higher amplitude (than at least the $r$-band). In both of these epochs, we can see short-term variability in the $g$-band, suggesting the presence of an additional component contributing to the modulation of the light curve.
To characterise the short term variability of GOTO 0650 we have computed the generalized Lomb-Scargle periodogram (LSP; Lomb, 1976; Scargle, 1982; VanderPlas, 2018) for each light curve as implemented in astropy version 6.1.2. In cases where our light curves contained no gaps, we also computed the power spectrum using the fast Fourier transform as implemented in NumPy version 2.0.1. We normalised our power spectra according to:
$P=\frac{2\delta t}{N}|DFT|^{2},$ (1)
where $\delta t$ is the time sampling of the light curve and $N$ is the number of points in the light curve. This normalisation is defined such that the integrated power spectrum yields the variance of the light curve (Vaughan et al., 2003). We normalised our LSPs according to:
$P={2\delta t}P_{\rm LSP},$ (2)
where $\delta t$ is the time sampling of the light curve and $P_{\rm LSP}$ represents the unnormalised LSP powers (equivalent to $|DFT|^{2}/N$ for uniformly sampled time series).
To estimate the confidence levels of the proposed signals in our power spectra, we model the power spectrum continuum analytically following the method described in Appendix A of Vaughan (2005). We compute confidence thresholds by scaling our power spectrum model by $-2{\rm ln}(\epsilon/n^{\prime})$, where $\epsilon$ is the desired false alarm probability and $n^{\prime}$ is the number of power spectrum points used to fit the model. The power spectra for each of the light curves is shown in the right panel of Figures 2 and 3. We found a periodicity of $\rm P_{\omega}=148.2$ s with more than $99.99\%$ confidence only in the $g$-band of epoch 4, though we also note a non-statistically significant signal at the same frequency in the r band of the same epoch. Unfortunately, the weather conditions during epoch 5 were far from ideal, with variable cloud coverage; this is reflected in our ability to obtain differential photometry, as seen in the corresponding epoch of Fig 3. In turn, we are not able to recover any signal from the power spectrum but, despite this, the $g$-band lightcurve from that same epoch exhibits short-term flares qualitatively similar to those observed in epoch 4. We derive the analytical false alarm probability of our epoch 4 $g$-band detection to be $1.16\times 10^{-6}$; the probability of detecting a signal of this significance in at least one of our 14 power spectra by chance (i.e., when there is no underlying signal) is 0.0016 per cent.
Figure 2: Relative light curves of all three cameras for epochs $1-3$ (left panels), along with their respective periodograms (right panels). The vertical shaded region in the power spectra indicates the frequency at which we detect a statistically significant signal in epoch 4.
Figure 3: Same as Fig. 2 but for epochs 4 and 5. For the $g$-band periodogram in epoch 4, the 99.99 per cent confidence threshold is shown with a dotted line.
To determine the centre and $1\sigma$ uncertainty of the frequency detected in epoch 4 we have performed a bootstrap analysis (e.g. Ivezić et al., 2014). In Figure 4, we present the bootstrapped frequency distributions as well as the LSPs for each band in epoch 4. For the LSPs, 99.99 per cent confidence thresholds are shown. These confidence thresholds were estimated by simulating 10,000 white-noise light curves for each camera. Each simulated light curve had the same time sampling as the observed light curve, and fluxes were generated from the following Gaussian $f_{i}\sim\mathcal{N}(\bar{f},\sigma_{f_{i}})$, where $\bar{f}$ represents the mean flux of the observed light curve and $\sigma_{f_{i}}$ represents the uncertainties on the observed flux values. As can be seen in Figure 3, the periodograms of Epoch 4 are white-noise-dominated beyond 5 mHz, and so our estimated confidence thresholds are unlikely to be significantly biased in this frequency range.
Figure 4: Left: frequency distributions inferred from bootstrapped Epoch 4 light curves. Right: Epoch 4 LSPs. In all panels, the dotted vertical line corresponds to a frequency of 6.7 mHz. The horizontal line represents the $99.99\%$ confidence threshold.
As can be seen in Figure 4, the 6.7 mHz signal is persistent in most bootstrapped light curves in the $g$-band. The 6.7 mHz signal is somewhat less persistent in the $r$-band, and even less so in the $i$-band. From these frequency distributions, we find peak frequencies of $6.73\pm 0.02$ mHz ($g$-band), $6.73\pm 0.02$ mHz ($r$-band), and $6.75\pm 0.02$ mHz (i-band). These frequencies correspond to periods of $148.5\pm 0.4$ s, $148.7\pm 0.4$ s, and $148.2\pm 0.5$ s, respectively.
To test the veracity of the signal, we have conducted power spectra of the other stars as well as the background itself, in case the signal was produced by artefacts from the electron noise in the detectors. We have not found any significant signal in any of these. In Figure 5, we present the LSP of the raw counts in the Epoch 4 $g$-band light curve for GOTO 0650 (top panel), along with the LSP of the corresponding local background (bottom panel). There is a strong peak in the periodogram of the raw flux at 6.7 mHz that is not seen in the periodogram of the local background. Since the periodogram in the top panel was computed from the raw flux, this figure shows that the 6.7 mHz signal cannot be attributed to our reference stars. Moreover, the lack of any prominent signal in the local background makes it unlikely that the 6.7 mHz signal is a result of some instrumental/systematic effect.
Figure 5: Top: LSP of GOTO 0650’s raw flux from Epoch 4 in the $g$-band; bottom: LSP of the corresponding local background.
4 Discussion
In Sec. 3, we have unambiguously found evidence of a periodic signal with a period of $148.5\pm 0.4$ s in epoch 4 during the dip in between the main outburst and the beginning of the rebrightenings/echo-outburst of the newly discovered WZ-Sge-type CV GOTO 0650. Unfortunately, the weather conditions did not allow us to recover the signal from epoch 5, but the light curve from this night looks qualitatively similar. But, what is the origin of the 148.5 seconds signal observed in GOTO 0650? In this section, we will briefly explore the different physical mechanisms that can produce the observed signal.
This signal is present in $g$-band with its amplitude decreasing toward longer wavelengths with a tentative detection in $r$-band. This suggests a blue spectral component generating the detected signal, similar to the one seen in IPs. In addition to this, the frequency of the observed signal is comparable with the dominant signal of a typical IP. This makes the GOTO 0650 a WZ Sge-type IP the most obvious interpretation, nonetheless we will consider other scenarios.
When there is more than one periodicity in a CV, typically, their periodograms exhibit an excess of power due to the interference of these signals, a.k.a. the beat period. We can use the presence of “regular” superhumps to estimate the orbital period. A period of $\rm P_{sh}=91.5$ min places GOTO 0650 below the period gap of CVs as its orbital period would be 10% shorter than $\rm P_{sh}$ (e.g. Kato et al., 2008). Assuming these values we can estimate the beat frequency as $\rm\nu_{beat}=\nu_{orb}-\nu_{sh}$ (where $\rm\nu_{beat},\ \nu_{orb}\ and\ \nu_{sh}$ are the beat, orbital and superhump frequency respectively), this results in a beat frequency $\rm\nu_{beat}\simeq 20~{}\mu Hz$, $\sim 300$ times slower than the signal we observe and more than a factor of $\sim 30$ from other WZ Sge-IPs (Woudt et al., 2012). Dwarf nova oscillations (DNOs) have been proposed to explain some variable periodicity in outbursting CVs (e.g. Marsh & Horne, 1998; Woudt & Warner, 2002), these are vertically extended regions of the disk or blobs being irradiated by the hot WD and boundary layer, which produce a signal corresponding to their Keplerian orbit. These signals are typically of the order of 10 seconds and evolve to slightly lower frequencies as the outburst evolves (up to 40s). We do not detect any significant signal during the main outburst, therefore we discard this scenario. The accretion-induced heating from the outburst makes non-radial pulsations of the WD also unlikely since the temperature of the WD should be too hot to be in the instability strip (c.f. Clemens, 1993; Toloza et al., 2016, but also see Szkody 2021), given its pre-outburst temperature (Killestein et al. submitted). Finally, Veresvarska et al. (2024) proposed a precessing inner disk producing QPOs in a handful of CVs, however, these observed frequencies are typically one order of magnitude slower. Therefore we conclude that GOTO 0650 is likely a WZ Sge-IP, but, the stability of the signal would need to be tested once the system is back to quiescent levels.
Magnetism has been invoked to explain several properties of WZ Sge stars (see Sect. 1). WZ Sge itself is probably the most remarkable example of this class. This source exhibits fast optical, UV and X-ray oscillations in quiescence (Robinson et al., 1978; Patterson, 1980; Patterson et al., 1998; Skidmore et al., 1999) around 27.87s that have been associated with the spin period of the WD (Patterson et al., 1998). This signal is absent during outbursts owing to the combination of low magnetic field and/or low mass accretion rate from the donor, however, this interpretation has been challenged (c.f. Knigge et al., 2002), leaving the origin of this signal an enduring enigma. CC Scl and ASASSN-18fk are also proposed to harbour a slowly spinning WD (Woudt et al., 2012; Pavlenko et al., 2019; Paice et al., 2024). Another remarkable example of an intermediate polar exhibiting super-outbursts is V455 And (Araujo-Betancor et al., 2005; Bloemen et al., 2013), this system exhibits a spin period of 67.6 seconds with the dominant signal in the power spectrum being twice the spin period; however, it does not exhibit echo outbursts like GOTO 0650 and WZ Sge. The timescale of the signal observed in GOTO 0650 falls in-between that of CC Scl and V455 And, and the absence of the signal during outburst could be attributed to the combination of low magnetic field and/or mass transfer rate similar to what has been proposed for WZ Sge (Patterson et al., 1998). We then propose the observed signal as being produced by the spin period (or its first harmonic).
Under this assumption, we can estimate the magnetic field assuming canonical values for typical WZ Sge stars from Knigge et al. (2011), i.e. $\dot{\rm M}_{tr}\simeq 10^{-11}~{}\mathrm{M_{\odot}\,yr^{-1}}$, and a typical $0.8M_{\odot}$ WD (Zorotovic et al., 2011; Pala et al., 2017). Assuming the magnetosphere of the white dwarf truncates the disk at the co-rotation radius, i.e. $\rm r_{M}=\xi r_{A}=r_{CO}$ (with $\rm\xi\simeq 0.5$ e.g. Ghosh & Lamb, 1979; Long et al., 2005), we can estimate the the magnetic field, $\rm B$, to be $\sim 1-5\times 10^{4}~{}{\rm G}$ for $\rm P_{spin}=1-2\times P_{\omega}$ and the $\dot{\rm M}_{acc}$ required to push the magnetosphere to the surface of the WD would be $\rm 6\times 10^{-9}~{}M_{\odot}~{}yr^{-1}$ and $\rm 3\times 10^{-8}~{}M_{\odot}~{}yr^{-1}$ for $\rm P_{\omega}=1$ and $2\times P_{spin}$ respectively. Given that we only detect significant periodicity at low luminosities $\rm L\lesssim 10\times L_{quiescence}$ and the accretion luminosity is proportional to the mass accretion rate (${\rm L}_{acc}\propto\dot{\rm M}_{acc}$), we suggest that $\rm P_{\omega}$ is the fundamental frequency of the spin period. If confirmed, GOTO 0650 would not only be another example of how magnetic fields can be dynamically important in highly evolved CVs but also would bridge the gap between intermediate- and fast- rotating AWDs.
Acknowledgements
We are grateful to many amateur astronomers for making their data publicly available in AAVSO. NCS and DLC acknowledges support from the Science and Technology Facilities Council (STFC) grant ST/X001121/1. ZAI, FV and DA acknowledges support STFC grants ST/V001000/1 and ST/X508767/1. AC acknowledges support from the Royal Society Newton International Fellowship grant NF170803. Based upon observations carried out at the Observatorio Astronómico Nacional on the Sierra San Pedro Mártir (OAN-SPM), Baja California, México.
Affiliations
1Department of Physics, University of Warwick, Gibbet Hill Road, Coventry, CV4 7AL, UK
2Department of Physics & Astronomy. University of Southampton, Southampton SO17 1BJ, UK.
5Instituto de Astronomía, Universidad Nacional Autónoma de México, Carretera Tijuana-Ensenada Km. 107, Ensenada, B.C. 22860, México
7Centre for Extragalactic Astronomy, Department of Physics, Durham University, South Road, Durham, DH1 3LE
10NAF – Osservatorio Astronomico di Capodimonte, Salita Moiariello 16, I-80131 Naples, Italy
8Advanced Research Center for Space Science and Technology, College of Science and Engineering, Kanazawa University, Kakuma, Kanazawa, Ishikawa 920-1192, Japan
9Cluster for Pioneering Research, Institute of Physical and Chemical Research (RIKEN), 2-1 Hirosawa, Wako, Saitama 351-0198, Japan
3South African Astronomical Observatory, PO Box 9, Observatory, 7935, Cape Town, South Africa
4Department of Astronomy, University of Cape Town, Private Bag X3, Rondebosch 7701, South Africa
6SUPA School of Physics and Astronomy, University of St Andrews, North Haugh, St Andrews KY16 9SS, Scotland, UK
Data Availability
The data underlying this article is publicly available in AAVSO https://www.aavso.org and ZTF https://www.ztf.caltech.edu/ztf-public-releases.html. The remaining data will be shared on reasonable request to the corresponding author.